# Probing nuclear liquid-gas phase transition with isospin correlations

### v1

@@ III Summary
We investigated the net-baryon number density fluctuations in the vicinity of the nuclear liquid-gas phase transition in the isospin-symmetric matter. To this end, we used the parity doublet model in the mean-field approximation to account for the critical behavior. In particular, we have studied the susceptibilities of the net-proton and net-neutron numbers, as well as their correlations up to the fourth order. These quantities show nonmonotonic structures at the phase boundary and are one of the main tools used in experimental searches for the QCD critical point.
@@
We discussed the effects of the correlations between protons and neutrons and found that they are non-trivial. Consequently, the net-proton number fluctuations have a qualitatively different structure than the net-baryon number fluctuations. Our results emphasize the importance of the interactions between baryons for higher-order fluctuations. In particular, they bring significant and nontrivial differences in the critical behavior of the net-proton and net-baryon number density fluctuations close to the critical endpoint of the liquid-gas phase transition. We hope they could be useful in searching for critical points in the QCD phase diagram and interpreting the critical properties of the matter created in heavy-ion collisions.
@@ Acknowledgments
This work was supported by the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science (M.M.). This work is partly funded by the Polish National Science Centre (NCN) under OPUS Grant No. 2022/45/B/ST2/01527 (C.S. and K.R.). K.R. also acknowledges the support of the Polish Ministry of Science and Higher Education. The work of C.S. was supported in part by the World Premier International Research Center Initiative (WPI) under MEXT, Japan.

### v2

@@ IV Summary
We investigated the net-baryon number density fluctuations close to the nuclear liquid-gas phase transition in the isospin-symmetric matter. To this end, we used the parity doublet model in the mean-field approximation to account for the critical behavior. In particular, we have studied the susceptibilities of the net-proton and net-neutron numbers, as well as their correlations up to the fourth order. These quantities show nonmonotonic structures at the phase boundary and are one of the main tools used in experimental searches for the QCD critical point. We discussed the effects of the correlations between protons and neutrons and found that they are non-trivial. Consequently, the net-proton number fluctuations have a qualitatively different structure than the net-baryon number fluctuations.
We have also studied the factorial cumulants. We find that the qualitative structure of the baryon-number factorial cumulant is preserved when the proton-number factorial cumulant is considered. Furthermore, the difference between baryon- and proton-number factorial cumulants is directly proportional to a linear combination of the proton-neutron correlation functions. We demonstrated that although it is nontrivial, it has a qualitatively similar structure to the factorial cumulants themselves. We expect these results to hold in the vicinity of the liquid-gas phase transition.
@@
Our results emphasize the importance of the interactions between baryons for higher-order fluctuations. In particular, they bring significant and nontrivial differences in the critical behavior of the net-proton and net-baryon number density fluctuations close to the critical endpoint of the liquid-gas phase transition. We hope they could be useful in searching for critical points in the QCD phase diagram and interpreting the critical properties of the matter produced in heavy-ion collisions.
@@ Acknowledgments
This work was supported by the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science (M.M.). This work is partly funded by the Polish National Science Centre (NCN) under OPUS Grant No. 2022/45/B/ST2/01527 (C.S. and K.R.). K.R. also acknowledges the support of the Polish Ministry of Science and Higher Education. The work of C.S. was supported in part by the World Premier International Research Center Initiative (WPI) under MEXT, Japan. We acknowledge interesting discussions with Volker Koch and Larry McLerran. We also thank the anonymous referee for their constructive comments and suggestion to extend the study to factorial cumulants.